# Spontaneous Speech-Based Suicide Risk Detection Using Whisper and Large Language Models

### v1

@@ 5.1 Comparison of Different Tuning and Fusion Strategies
| 1 | Whisper | APFT | 0.758 0.029 | |||
| 2 | Whisper | PEFT | 0.633 0.036 | |||
| 3 | Baichuan2 | APFT | 0.763 0.015 | |||
| 4 | Baichuan2 | PEFT | 0.698 0.043 | |||
| 5 | Qwen1.5 | APFT | 0.675 0.034 | |||
| 6 | Qwen1.5 | PEFT | 0.737 0.024 | |||
| 7 | Whisper | APFT | Baichuan2 | APFT | CC | 0.768 0.021 |
| 8 | Whisper | PEFT | Baichuan2 | PEFT | CC | 0.718 0.040 |
| 9 | Whisper | APFT | Qwen1.5 | APFT | CC | 0.710 0.029 |
| 10 | Whisper | PEFT | Qwen1.5 | PEFT | CC | 0.752 0.014 |
| 11 | Whisper | APFT | Qwen1.5 | APFT | IC | 0.730 0.033 |
| 12 | Whisper | APFT | Qwen1.5 | PEFT | IC | 0.743 0.026 |
| 13 | Whisper | PEFT | Qwen1.5 | PEFT | IC | 0.694 0.019 |

### v2

@@ 5.1 Comparison of Different Tuning and Fusion Strategies
1 | Whisper | APFT | 0.758 0.029 | |||
2 | Whisper | PEFT | 0.633 0.036 | |||
3 | Baichuan2 | APFT | 0.763 0.015 | |||
4 | Baichuan2 | PEFT | 0.698 0.043 | |||
5 | Qwen1.5 | APFT | 0.675 0.034 | |||
6 | Qwen1.5 | PEFT | 0.737 0.024 | |||
7 | Whisper | APFT | Baichuan2 | APFT | CC | 0.768 0.021 |
8 | Whisper | PEFT | Baichuan2 | PEFT | CC | 0.718 0.040 |
9 | Whisper | APFT | Qwen1.5 | APFT | CC | 0.710 0.029 |
10 | Whisper | PEFT | Qwen1.5 | PEFT | CC | 0.752 0.014 |
11 | Whisper | APFT | Qwen1.5 | APFT | IC | 0.730 0.033 |
12 | Whisper | APFT | Qwen1.5 | PEFT | IC | 0.743 0.026 |
13 | Whisper | PEFT | Qwen1.5 | PEFT | IC | 0.694 0.019 |